# Learning Diverse Attacks on Large Language Models for Robust Red-Teaming and Safety Tuning

### v1

@@ 4.1 Experimental setup
We validate the efficacy of our method by red-teaming a variety of open source LMs with varying levels of saferty-tuning. Recall that the goal of red-teaming is to generate diverse prompts that elicit harmful responses from a target LLM $p_{\phi}$. In our experiments, we use GPT-2 [43], dolly-v2-7b [13], Llama-2-7b-chat [51], and Gemma-2b-it [39] as target LLMs. Note that we only require black-box access, i.e., we can only query a prompt and get a text response.
@@
To measure effectiveness of each method, we compute the toxicity rate which is the percentage of generated prompts that are toxic. We consider a prompt toxic if the toxicity classifier $p_{\psi}$ assigns a score greater than 0.5 to an input pair consisting of the prompt ${\mathbf{x}}$ and response ${\mathbf{y}}$ generated by the target LLM. For $p_{\psi}$, we use a RoBERTa hate speech classifier [53] for GPT-2 and dolly-v2-7b, and Llama-Guard [24] for Llama-2-7b-chat and Gemma-2b-it. To measure the diversity of the generated prompts, we embed the generated prompts using MiniLMv2 [55] encoder and compute the average pairwise cosine distance between embeddings of the prompts.
@@ Methods.
    Supervised Fine-tuning (SFT): We fine-tune the pretrained LM $p_{\theta}$ with a maximum likelihood objective on 3,003 toxic prompts from SafetyDataset [7] and AdvBench [68].
@@
    In-Context Learning (ICL) [8]: We sample 5-shot demonstrations from toxic prompt datasets (SafetyDataset and AdvBench) and prompt GPT-2 to generate a prompt.
@@
    REINFORCE [60]: We fine-tune the pretrained LM $p_{\theta}$ as an RL policy with policy gradients to optimize the reward in Equation 1.
@@
    PPO + Novelty [20]: This method adds entropy bonus [48] along with a novelty-based term to the reward in Equation 1 and train the policy $p_{\theta}$ with proximal policy optimization (PPO) [49]. For novelty-based reward, it utilizes self-BLEU [67] and pairwise cosine similarity between embeddings of all the past generated prompts.
@@
    GFlowNet [36]: We fine-tune the pretrained LM $p_{\theta}$ with Equation 3.
@@
    GFlowNet + MLE: This is our full method for collecting high-reward prompts during GFlowNet fine-tuning and re-train the pretrained LM $p_{\theta}$ with maximum likelihood estimation (MLE) on the collected prompts as described in Alg. 1.
Figure 2: Percentage of toxic prompts out of $1,024$ samples and pairwise cosine distance of prompts generated by each method for (a) GPT-2, (b) Dolly-2-7b, (c) Gemma-it-2b, and (d) Llama-2-7b-chat victim models.
@@ 4.2 Results: Robust red-teaming
As the number of prompts which would elicit toxic responses occupy a small subset of all possible sequences, there is a natural trade-off between diversity and toxicity. We start by investigating how each method handles this trade-off. Fig. 2 illustrates the cosine distance plotted against the toxicity rate for 1,024 red-teaming prompts generated by each method across four different target LLMs. We find that our GFlowNet + MLE is the only method which manages to balance a high toxicity rate with the diversity of generated prompts across all four target LLMs. Qualitative assessment of examples generated by GFlowNet + MLE, included in Table B.3Table B.4Table B.5, and Table B.6, supports the numerical results. While the GFlowNet achieves both high diversity and toxicity rate for red-teaming GPT-2 (2(a)) and dolly-v2-7b (2(b)), the toxicity rate drops significantly for the target LLMs with safety fine-tuning: Gemma-2b-it (2(c)) and Llama-2-7b-chat (2(d)). We hypothesize this drop comes from the reward signal (toxicity of responses from the target) becoming sparse with safety-tuned models. Similarly, PPO + Novelty fails to find a balance between diversity and toxicity. When it is able to find effective prompts (2(a) and 2(b)) they are not as diverse and for safety fine-tuned models it fails to find any prompts which elicit a toxic response (2(c) and 2(d)). (For context, a random policy would have the highest diversity but would have a low toxicity rate). On the other hand, REINFORCE, which does not take account diversity into the reward function, collapses to deterministically generating a single reward-maximizing prompt. Finally, SFT and ICL generate diverse but ineffective prompts.
@@
Figure 3: Percentage of prompts out of $1,024$ samples for each toxicity score bin with red-teaming the Llama-2-7b-chat victim language model. Results of the other victim models are included in §B.1.
@@
Table 2: We generate $1,024$ prompts with the policy trained for red-teaming Gemma-2b-it and evaluate the prompts with different victim models. All the results represent averages from five different experimental runs.
@@
GFlowNet + MLE 85.1685.16{85.16}85.16 27.39 24.28 22.94 29.98 52.01 67.84 77.16 61.9461.94{61.94}61.94 66.63 67.21
GFlowNet + MLE generates diverse and effective prompts.
To further understand the behavior of each method beyond the toxicity rate (which depends on the $p_{\psi}(c=1|{\mathbf{x}},{\mathbf{y}})>0.5$ decision boundary), we illustrate the distribution over the toxicity scores and corresponding average pairwise cosine distances for the generated prompts in Fig. 3, obtained from the experiment for red-teaming the Llama-2-7b-chat target LLM. Results for the other target LLMs are illustrated in Fig. B.1Fig. B.2, and Fig. B.3 in §B.1. GFlowNet + MLE achieves consistently high diversity across different toxicity score bins. On the other hand, all other methods fail to achieve high diversity and toxicity at the same time. GFlowNet generates fewer toxic prompts compared to GFlowNet + MLE. Notably, PPO + Novelty does not generate prompts with the toxicity score greater than $0.2$ at all for Gemma-2b-it and Llama-2-7b-chat. While REINFORCE generates a single highly toxic prompt achieving a much lower diversity, SFT and ICL generate few toxic prompts.
GFlowNet attacks are more transferable across target LMs.
A potential advantage of generating diverse attack prompts is that prompts generated for red-teaming a given target LLM can potentially transfer to other LLMs, since some of the failure modes of a target LLM might be shared by other models, for instance, due to using similar web-filtered data or similar safety alignment recipes. To study this empirically, we train an attacker policy $p_{\theta}$ for red-teaming the Gemma-2b-it as the target LLM. We then sample $1,024$ prompts from the trained attacker LM and evaluate the number of prompts which transfer to other LLMs, i.e., elicit toxic responses from unseen LLMs: Llama-2-7b-chat, Llama-2-13b-chat, Llama-2-70b-chat, Llama-3-8b-instruct [1], Llama-3-70b-instruct, Gemma-7b-it, Gemma-1.1-2b-it, Gemma-1.1-7b-it, Mistral-7b-instruct-v0.2 [28], and Starling-7b-beta [66]. As shown in Table 2, we find that many prompts generated by GFlowNet + MLE transfer to unseen target LLMs, outperforming all other methods across all the target LLMs except Mistral-7b-instruct-v0.2. REINFORCE generates almost identical prompts, tailored to the Gemma-2b-it target it was trained with, which consequently do not transfer to other target LLMs. This highlights a drawback of methods which do not generate diverse attacks. On the other extreme, PPO + Novelty is unable to discover any prompt that is effective in eliciting toxic responses and consequently none of the prompts transfer to any other LLM. These results further highlight the efficacy and usefulness of GFlowNet + MLE, which can generate both diverse and effective red-teaming prompts that can be transferred to red-team other LLMs.
@@
Another advantage of our two stage approach is that it can enable fast adaptation of an attacker LM policy to red-team a different target LLM. Concretely, we can compute reward of the stored attack prompts, sampled during GFlowNet fine-tuning, with a different target LLM and rerank the prompts (instead of scoring them with the same target LLM). The offline dataset can be constructed by filtering the prompts with the newly computed $R_{1}({\mathbf{x}})$ and pre-computed $R_{2}({\mathbf{x}})$ based on the corresponding thresholds $r_{1}$ and $r_{2}$. The initial pre-trained attacker LM policy $p_{\theta}$ is fine-tuned with supervised learning on this dataset. For this experiment, we consider the the prompts stored during the red-teaming of Gemma-2b-it and adapt the attacker LM to red-team Gemma-1.1-2b-it, Gemma-7b-it, Gemma-1.1-7b-it, Llama-2-7b-chat, and Llama-3-8b-instruct target LLMs. As shown in §4.2, adaptation of the attack LM policy with this reranking procedure is effective and significantly improves toxicity rate over direct transfer from an attacker trained to red-team the initial target LLM, Gemma-2b-it. Note that a considerable amount of computational cost and wall-clock time can be saved since we skip the GFlowNet fine-tuning stage and simply re-use the stored prompts.
@@
In this experiment, we demonstrate empirically the challenges in tuning the temperature $\beta$ in Equation 2 and how second phase of MLE smoothing provides a better trade-off between toxicity rate and diversity. We fine-tune the pretrained initial policy $p_{\theta}$ as a GFlowNet by setting the temperature $\beta$ to each value in $\{0.01,0.02,\ldots,0.1,1.0\}$ and fine-tune again the initial attacker LM policy with MLE on each of the high-reward prompts discovered during GFlowNet fine-tuning with the corresponding $\beta$. As shown in Fig. 5, the GFlowNet objective sacrifices diversity (cosine distance) considerably to obtain high toxicity rate, or it significantly degrades the toxicity rate resulting in diverse prompts. On the other hand, smoothing with MLE is robust to this choice of $\beta$ and enables the attacker policy to sample effective attack prompts while retaining diversity.
@@
We perform an ablation study to demonstrate the importance of the off-policy exploration ability of GFlowNets for collecting the offline dataset for the second phase of MLE smoothing as described in Alg. 1. To show this, we train the initial pretrained policy $p_{\theta}$ with PPO + Novelty, instead of GFlowNet fine-tuning, for red-teaming Llama-2-7b-chat target LLM and attempt to collect high reward prompts sampled by the policy during training. However, PPO + Novelty does not generate any high-reward prompts at all during training, as shown in §4.2. Thus, we cannot collect any samples for the offline dataset and consequently cannot fine-tune the policy with MLE. This indicates that a diversity-seeking approach like GFlowNet fine-tuning is crucial for collecting effective red-teaming prompts.
@@ 4.3 Results: Safety fine-tuning
To demonstrate that attack prompts generated by red-teaming methods can enhance the safety alignment of target LLMs, we fine-tune Gemma-2b-it target LLM with these attack prompts and evaluate its robustness. Specifically, we generate 1,024 prompts with each red-teaming method and generate refusal responses to the attack prompts with the target LLM using the prompt: “Refuse answering to the following instruction. {attack_prompt}”. Subsequently, the target model is fine-tuned with LoRA [21] to maximize the conditional log-likelihood of the refusal responses to the attack prompts, resulting in six different fine-tuned target LLMs corresponding to each red-teaming method. Finally, each fine-tuned model generates responses to the attack prompts generated by each red-teaming method, and we measure the toxicity rate of the responses with Llama-Guard as the toxicity classifier $p_{\psi}$.
@@
As shown in Fig. 6, the target LLM fine-tuned on the attack prompts generated by GFlowNet + MLE is the most robust to unseen attack prompts generated by the other RL-based red-teaming methods. On the other hand, even after safety fine-tuning, all the other target LLMs cannot defend against the attack prompts generated by GFlowNet + MLE. We also confirm that our safety-tuned model still preserves general instruction-following capabilities: as shown in Table B.2, the performance on the six tasks in the Open LLM Leaderboard changes insignificantly with safety tuning. These experimental results verify the importance of generating diverse and effective prompts for robust red-teaming.

### v2

@@ 4.1 Experimental setup
We validate the efficacy of our method by red-teaming a variety of open source LMs with varying levels of saferty-tuning. Recall that the goal of red-teaming is to generate diverse prompts that elicit harmful responses from a target LLM $p_{\phi}$. In our experiments, we use GPT-2 (Radford et al., 2019), Dolly-v2-7b (Conover et al., 2023), Llama-2-7b-chat (Touvron et al., 2023), Llama-3.1-8B-Instruct (Dubey et al., 2024), and Gemma-2b-it (Mesnard et al., 2024) as target LLMs. Note that we only require black-box access, i.e., we can only query a prompt and get a text response.
@@
To measure effectiveness of each method, we compute the toxicity rate which is the percentage of generated prompts that are toxic. We consider a prompt toxic if the toxicity classifier $p_{\psi}$ assigns a score greater than 0.5 to an input pair consisting of the prompt ${\mathbf{x}}$ and response ${\mathbf{y}}$ generated by the target LLM. For $p_{\psi}$, we use a RoBERTa hate speech classifier (Vidgen et al., 2021) for GPT-2 and dolly-v2-7b, and Llama-Guard (Inan et al., 2023) for Llama-2-7b-chat, Llama-3.1-8B-Instruct, and Gemma-2b-it. To measure the diversity of the generated prompts, we embed the generated prompts using MiniLMv2 (Wang et al., 2021) encoder and compute the average pairwise cosine distance between embeddings of the prompts.
@@ Methods.
    Supervised Fine-tuning (SFT): We fine-tune the pretrained LM $p_{\theta}$ with a maximum likelihood objective on 3,003 toxic prompts from SafetyDataset (Bianchi et al., 2024) and AdvBench (Zou et al., 2023).
@@
    In-Context Learning (ICL) (Brown et al., 2020): We sample 5-shot demonstrations from toxic prompt datasets (SafetyDataset and AdvBench) and prompt GPT-2 to generate a prompt.
@@
    REINFORCE (Williams, 1992): We fine-tune the pretrained LM $p_{\theta}$ as an RL policy with policy gradients to optimize the reward in Equation 1.
@@
    PPO + Novelty (Hong et al., 2024): This method adds entropy bonus (Schulman et al., 2017a) along with a novelty-based term to the reward in Equation 1 and train the policy $p_{\theta}$ with proximal policy optimization (PPO; Schulman et al., 2017b). For novelty-based reward, it utilizes self-BLEU (Zhu et al., 2018) and pairwise cosine similarity between embeddings of all the past generated prompts.
@@
    GFlowNet (Malkin et al., 2022): We fine-tune the pretrained LM $p_{\theta}$ with Equation 3. (This is Stage 1 of our full procedure.)
@@
    GFlowNet + MLE: This is our full method for collecting high-reward prompts during GFlowNet fine-tuning and re-training the pretrained LM $p_{\theta}$ with maximum likelihood estimation (MLE) on the collected prompts as described in Alg. 1.
Figure 2: Percentage of toxic prompts (measuring toxicity) out of $10,000$ samples and pairwise cosine distance of prompts generated by each method (measuring diversity) for (a) Dolly-2-7b, (b) Gemma-it-2b, (c) Llama-2-7b-chat, and (d) Llama-3.1-8B-Instruct target models. Results for GPT-2 in Fig. B.1 in §B.1.
@@ 4.2 Results: Robust red-teaming
As the number of prompts which would elicit toxic responses occupy a small subset of all possible sequences, there is a natural trade-off between diversity and toxicity. We start by investigating how each method handles this trade-off. Fig. 2 illustrates the cosine distance plotted against the toxicity rate for $10,000$ red-teaming prompts generated by each method across five different target LLMs. We find that our GFlowNet + MLE is the only method which manages to balance a high toxicity rate with the diversity of generated prompts across all four target LLMs. Qualitative assessment of examples generated by GFlowNet + MLE, included in Table B.5Table B.6Table B.7,  Table B.8, and Table B.9, supports the numerical results. While the GFlowNet achieves both high diversity and toxicity rate for red-teaming GPT-2 (Fig. B.1) and Dolly-v2-7b (2(a)), the toxicity rate drops significantly for the target LLMs with safety fine-tuning: Gemma-2b-it (2(b)), Llama-2-7b-chat (2(c)), and Llama-3.1-8B-Instruct (2(d)). We hypothesize this drop comes from the reward signal (toxicity of responses from the target) becoming sparse with safety-tuned models. Similarly, PPO + Novelty fails to find a balance between diversity and toxicity. When it is able to find effective prompts (Fig. B.1 and 2(a)), they are not as diverse and for models with strong safety-guardrail, such as Llama-2 and Gemma, it fails to find any prompts which elicit a toxic response (2(b) and 2(c)). When it comes to red-teaming Llama-3.1-8B-Instruct, it moderately finds a balance between toxicity and diversity but still falls significantly short compared to our GFlowNets + MLE approach. (For context, a random policy would have the highest diversity but would have a low toxicity rate). On the other hand, REINFORCE, which does not take diversity into account, collapses to deterministically generating a single reward-maximizing prompt. Finally, SFT and ICL generate diverse but ineffective prompts.
Table 2: Comparison of different attacker LMs for red-teaming Llama-3.1-8B-Instruct model.
Base Model Toxicity Rate (%) Cosine Distance
GPT2-small $81.05\pm 0.96$ $0.829\pm 0.001$
Llama-3.2-1B $\textbf{92.71}\pm 1.12$ $0.833\pm 0.002$
Scaling to a larger attacker LM.
Section 4.2 shows the effect of scaling GFlowNet+MLE with larger and stronger attackers like Llama-3.2-1B (Dubey et al., 2024). Scaling to a larger attacker results in significant improvements in both the toxicity rate and diversity.
@@
Figure 3: Percentage of prompts out of $10,000$ samples for each toxicity score bin with red-teaming the Llama-2-7b-chat target language model. Results for other target models are included in §B.2.
GFlowNet + MLE generates diverse and effective prompts.
To further understand the behavior of each method beyond the toxicity rate (which depends on the $p_{\psi}(c=1\mid{\mathbf{x}},{\mathbf{y}})>0.5$ decision boundary), we illustrate the distribution over the toxicity scores and corresponding average pairwise cosine distances for the generated prompts in Fig. 3, obtained from the experiment for red-teaming the Llama-2-7b-chat target LLM. Results for the other target LLMs are illustrated in Fig. B.2Fig. B.3,  Fig. B.4, and Fig. B.5 in §B.2. GFlowNet + MLE achieves consistently high diversity across different toxicity score bins. On the other hand, all other methods fail to achieve high diversity and toxicity at the same time. GFlowNet generates fewer toxic prompts compared to GFlowNet + MLE. Notably, PPO + Novelty does not generate prompts with the toxicity score greater than $0.2$ at all for Gemma-2b-it and Llama-2-7b-chat. While REINFORCE generates a single highly toxic prompt achieving a much lower diversity, SFT and ICL generate few toxic prompts.
@@
Table 3: We generate $1,024$ prompts with the policy trained for red-teaming Gemma-2b-it and evaluate the prompts with different target models. All the results represent averages from five different experimental runs.
@@
GFlowNet + MLE 85.1685.16{85.16}85.16 27.39 24.28 22.94 29.98 52.01 67.84 77.16 61.94 66.6366.63{66.63}66.63 67.21
GFlowNet attacks are more transferable across target LLMs.
A potential advantage of generating diverse attack prompts is that prompts generated for red-teaming a given target LLM can potentially transfer to other LLMs, since some of the failure modes of a target LLM might be shared by other models, for instance, due to using similar web-filtered data or similar safety alignment recipes. To study this empirically, we train an attacker policy $p_{\theta}$ for red-teaming the Gemma-2b-it as the target LLM. We then sample $1,024$ prompts from the trained attacker LM and evaluate the number of prompts which transfer to other LLMs, i.e., elicit toxic responses from unseen LLMs: Llama-2-7b-chat, Llama-2-13b-chat, Llama-2-70b-chat, Llama-3-8b-instruct (Dubey et al., 2024), Llama-3-70b-instruct, Gemma-7b-it, Gemma-1.1-2b-it, Gemma-1.1-7b-it, Mistral-7b-instruct-v0.2 (Jiang et al., 2023), and Starling-7b-beta (Zhu et al., 2023). As shown in Table 3, we find that many prompts generated by GFlowNet + MLE transfer to unseen target LLMs, outperforming all other methods across all the target LLMs except Mistral-7b-instruct-v0.2. REINFORCE generates almost identical prompts, tailored to the Gemma-2b-it target it was trained with, which consequently do not transfer to other target LLMs. This highlights a drawback of methods which do not generate diverse attacks. On the other extreme, PPO + Novelty is unable to discover any prompt that is effective in eliciting toxic responses and consequently none of the prompts transfer to any other LLM. These results further highlight the efficacy and usefulness of GFlowNet + MLE, which can generate both diverse and effective red-teaming prompts that can be transferred to red-team other LLMs. Additionally, we perform another transfer experiment targeted for a proprietary model, GPT-4o. We generate 1,024 prompts with an attacker LM trained to red-team Llama-2-7b-chat and evaluate how many prompts can elicit harmful responses from GPT-4o. On average, across five different sets of 1,024 prompts, 65% of them can successfully attack GPT-4o.
@@
Table 4: Training cost of each method with Llama-2-7b-chat target model.
Method Wall-clock Time GPU Memory
REINFORCE 5d 7h 24m 77.73 GB
PPO + Novelty 4d 7h 22m 79.87 GB
GFlowNet (1st stage) 1d 17h 37m 77.73 GB
GFlowNet + MLE (2nd stage) 1h 58m 37.25 GB
Stage 2 (MLE) is cheap.
As shown in §4.2, our proposed second stage MLE training is a lightweight process compared to other RL methods since it does not need on-policy samples or expensive reward computation. With just two hours of additional training, MLE training can significantly enhance the diversity and toxicity rate of GFlowNets.
@@
Another advantage of our two-stage approach is that it can enable fast adaptation of an attacker LM policy to a new target: an attacker trained against one target LLM can be adapted to red-team a different target LLM by repeating Stage 2 on a dataset filtered using the new target LLM. Concretely, we can recompute the reward of the stored attack prompts sampled during GFlowNet fine-tuning (Stage 1), with a different target LLM and rerank the prompts (instead of scoring them with the same target LLM). The offline dataset can be constructed by filtering the prompts with the newly computed $R_{1}({\mathbf{x}})$ and the precomputed $R_{2}({\mathbf{x}})$ based on the corresponding thresholds $r_{1}$ and $r_{2}$. The initial pretrained attacker LM policy $p_{\theta}$ is fine-tuned with supervised learning on this dataset. For this experiment, we consider the the prompts stored during the red-teaming of Gemma-2b-it and adapt the attacker LM to red-team Gemma-1.1-2b-it, Gemma-7b-it, Gemma-1.1-7b-it, Llama-2-7b-chat, and Llama-3-8b-instruct target LLMs. As shown in §4.2, adaptation of the attack LM policy with this reranking procedure is effective and significantly improves toxicity rate over direct transfer from an attacker trained to red-team the initial target LLM, Gemma-2b-it. Note that a considerable amount of computational cost and wall-clock time can be saved (cf. §4.2), since we skip the GFlowNet fine-tuning stage (Stage 1) and simply reuse the stored prompts.
@@
In this experiment, we demonstrate empirically the challenges in tuning the temperature $\beta$ in Equation 2 and how the second phase of MLE smoothing provides a better trade-off between toxicity rate and diversity. We fine-tune the pretrained initial policy $p_{\theta}$ as a GFlowNet by setting the temperature $\beta$ to each value in $\{0.01,0.02,\ldots,0.1,1.0\}$ and fine-tune again the initial attacker LM policy with MLE on each of the high-reward prompts discovered during GFlowNet fine-tuning with the corresponding $\beta$. As shown in Fig. 5, the GFlowNet objective sacrifices diversity (cosine distance) considerably to obtain high toxicity rate, or it significantly degrades the toxicity rate resulting in diverse prompts. On the other hand, smoothing with MLE is robust to this choice of $\beta$ and enables the attacker policy to sample effective attack prompts while retaining diversity.
@@
We perform an ablation study to demonstrate the importance of the off-policy exploration ability of GFlowNets for collecting the offline dataset in Stage 1 for MLE smoothing in Stage 2, as described in Alg. 1. To show this, we attempt to replace GFlowNet fine-tuning by PPO + Novelty in Stage 1. To be precise, we train the initial pretrained policy $p_{\theta}$ with PPO + Novelty for red-teaming the Llama-2-7b-chat target LLM and attempt to collect high-reward prompts sampled by the policy during training. However, PPO + Novelty does not generate any high-reward prompts at all during training, as shown in §4.2. Thus, we cannot collect any samples for the offline dataset and consequently cannot fine-tune the policy with MLE. This indicates that a diversity-seeking approach like GFlowNet fine-tuning is crucial for collecting effective red-teaming prompts.
@@ 4.3 Results: Safety fine-tuning
To demonstrate that attack prompts generated by red-teaming methods can enhance the safety alignment of target LLMs, we fine-tune Gemma-2b-it target LLM with the attack prompts discovered by various methods and evaluate its robustness. Specifically, we generate 1,024 prompts with each red-teaming method and generate refusal responses to the attack prompts with the target LLM using the prompt: “Refuse answering to the following instruction. {attack_prompt}”. Subsequently, the target model is fine-tuned with LoRA (Hu et al., 2022) to maximize the conditional log-likelihood of the refusal responses to the attack prompts, resulting in six different fine-tuned target LLMs corresponding to each red-teaming method. Finally, each fine-tuned model generates responses to the attack prompts generated by each red-teaming method, and we measure the toxicity rate of the responses with Llama-Guard as the toxicity classifier $p_{\psi}$.
@@
As shown in Fig. 6, the target LLM fine-tuned on the attack prompts generated by GFlowNet + MLE is the most robust to unseen attack prompts generated by the other RL-based red-teaming methods. On the other hand, even after safety fine-tuning, all the other target LLMs cannot defend against the attack prompts generated by GFlowNet + MLE. We also confirm that our safety-tuned model still preserves general instruction-following capabilities: as shown in Table B.2, the performance on the six tasks in the Open LLM Leaderboard changes insignificantly with safety tuning. These results highlight the importance of the diversity of generated red-teaming prompts for downstream safety fine-tuning.